12.       Suppose an economy's production function is specified by y = A(√K,√L), where A represents the productivity of resources, K the quantity of capital resources, and L the quantity of labor resources.

            A. Find potential output when A = 10, K = 49, and L =9.

            B. Find potential output when A = 10, K = 49, and labor resources are 3, 6, and 12.

            C. Plot the levels of output associated with 3, 6, 9, and 12 labor units; label the production function P'. What happens to this production function when productivity increases (the value of A is larger) and/or there is more capital?
Answer #1

A. Potential output when A = 10, K = 49, and L =9.

Y = A(√K.√L)

put the values

Y= 10 (√49.√9)

Y= 10. 7. 3

Y= 210

B. Potential output when A = 10, K = 49, and labor resources are 3, 6, and 12.

Y = A(√K.√L)

When L=3

Y = 10 (√49.√3)

Y= 10(7)(1.73)

Y= 121.1

When L=6

Y = 10 (√49.√6)

Y= 10(7)(2.44)

Y= 170.8

When L=9

Y= 210 (Solved in part A)

C) (12, 242 48) (9,210) Output (6, 170,8) (3, 121.1 (0, 0) 13 LabourThis Red Line is Production Function P'.

When productivity increases (the value of A is larger) and/or there is more capital the production Function shifts up. Because at the same level of capital and labour the productivity has increased.